OpenStudy (anonymous):

A baseball team plays in a stadium that holds 52000 spectators. With the ticket price at $12 the average attendance has been 19000. When the price dropped to $9, the average attendance rose to 26000. Assume that attendance is linearly related to ticket price. What ticket price would maximize revenue?

OpenStudy (anonymous):

P= price x= attendance

OpenStudy (anonymous):

x = kP +c for constants k and c

OpenStudy (anonymous):

when P=12, x= 19000

OpenStudy (anonymous):

when P= 9 , x= 26000

OpenStudy (anonymous):

now you can form two eqns, solve for the constants c and k

OpenStudy (anonymous):

19000 =12k +c 26000= 9k +c subtract the first from the second 7000= -3k --> k = -7000/3

OpenStudy (anonymous):

sub that into one of the other eqns say the first , so c= 19000-12k = 47000

OpenStudy (anonymous):

thats what ive done so far but nothing from there is making sense. I also have to describe what im doing in words so i prove i understand it.

OpenStudy (anonymous):

so x= (-7000/3) P +47000

OpenStudy (anonymous):

revenue = R = attendance times cost price

OpenStudy (anonymous):

so R = P [ (-7000/3)P +47000 ] differentiate with respect to P , we want to maximise R while varying P

OpenStudy (anonymous):

R = (-7000/3) P^2 + 47000P dR/dP = (-14000/3)P +47000 =0 for min/max

OpenStudy (anonymous):

P = 3

OpenStudy (anonymous):

now check, second derivative is d^2R/dP^2 = (-14000/3) which is less than zero , so we do get a maximum . so ticket price = $3 maximises revenue

OpenStudy (anonymous):

wit its wrong

OpenStudy (anonymous):

should be P = (-47000) / ( -14000/3 ) = $10.07 , that maximises revenue

OpenStudy (anonymous):

$10.07 is the right answer. I still however have no idea how you got it. We haven't went over any problems by differentiating. I have no idea what that is.

OpenStudy (anonymous):

well , this can be done by other methods , I am assuming you have done quadratic equations before , ie , you know how to find the vertex of a parabola?

OpenStudy (anonymous):

and you know that the vertex ( which is a maximum or a minimum ) occurs on the axis of symmetry , and axis of symmetry of y= ax^2 +bx +c is x=-b / (2a)

OpenStudy (anonymous):

yes that's what we're doing now

OpenStudy (anonymous):

how did I know ;)

OpenStudy (anonymous):

so if you go back to this eqn R = (-7000/3) P^2 + 47000P and find the axis of symmetry , which is P = -b/ (2a) , this will give you the x coordinate of the maximum point

OpenStudy (anonymous):

which is 10.07, the answer

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Ok i know how you got this answer but why is there no c in the ax^2+bx+c equation? and how did you get 2 P's in the equation R= (-7000/3)P^2 +47000P ?

OpenStudy (anonymous):

"why is there no c? " , there just isnt "where did the two Ps come from" , look at how revenue was calculated, revenue = price x number of people